### v1

# Differential Gated Self-Attention

Abstract
Transformers excel across a large variety of tasks but remain susceptible to corrupted inputs, since standard self-attention treats all query-key interactions uniformly. Inspired by lateral inhibition in biological neural circuits and building on the recent Differential Transformer’s use of two parallel softmax subtraction for noise cancellation, we propose Multihead Differential Gated Self-Attention (M-DGSA) that learns per-head input-dependent gating to dynamically suppress attention noise. Each head splits into excitatory and inhibitory branches whose dual softmax maps are fused by a sigmoid gate predicted from the token embedding, yielding a context-aware contrast enhancement. M-DGSA integrates seamlessly into existing Transformer stacks with minimal computational overhead. We evaluate on both vision and language benchmarks, demonstrating consistent robustness gains over vanilla Transformer, Vision Transformer, and Differential Transformer baselines. Our contributions are (i) a novel input-dependent gating mechanism for self-attention grounded in lateral inhibition, (ii) a principled synthesis of biological contrast-enhancement and self-attention theory, and (iii) comprehensive experiments demonstrating noise resilience and cross-domain applicability.
footnotetext: Code available at https://github.com/elyalyg/DGSA
1 Introduction
Attention-based Transformers (Vaswani et al., 2017) have revolutionized machine learning, driving breakthroughs in language, vision, and beyond. However, their uniform weighting of all query-key scores makes them vulnerable to input corruption - sensor noise in images or spurious tokens in text - which can be propagated and even amplified degrading performance in real-world datasets.
Existing techniques for enhancing attention robustness or reducing its complexity, typically either apply global regularization (e.g. dropout, weight decay) or rework the attention computation itself via sparsity or low-rank factorization (Wang et al., 2020; Choromanski et al., 2020; Zaheer et al., 2020). More recently, the Differential Transformer (DIFF Transformer) (Ye et al., 2024) introduces a head-wise subtraction of two paired softmax maps to cancel common-mode noise, but its use of a single learned scalar limits its ability to adapt to low-level noise patterns that vary at the granularity of individual tokens.
In this work, we address the challenge of noise-cancellation in self-attention by introducing Multihead Differential Gated Self-Attention (M-DGSA), a lightweight module that learns an input-dependent inhibitory gating per-head to dynamically adapt to and suppress attention noise.
Inspired by lateral inhibition in biological neural circuits, where neurons suppress neighboring activity to sharpen responses (Hartline, 1940; Kuffler, 1953), M-DGSA splits each head into excitatory and inhibitory branches. Dual softmax maps are fused via a sigmoid gate predicted from the token embedding, enabling fine-grained contrast enhancement tailored to each query-key pair. M-DGSA integrates seamlessly into off-the-shelf Transformer-based stacks with negligible overhead.
We evaluate M-DGSA on both vision and language benchmarks - namely CIFAR-10, CIFAR-100, FashionMNIST, and SVHN for vision, Rotten Tomatoes, IMDB, AGNews, and 20Newsgroups for language - under both clean and noisy settings. Empirically, M-DGSA consistently improves accuracy and noise resilience across these domains, outperforming vanilla Transformer (Vaswani et al., 2017), Vision Transformer (ViT) (Dosovitskiy et al., 2020), and Differential Transformer (Ye et al., 2024) baselines. These results underscore the practical benefit of adaptive, input-dependent inhibition for attention-based architectures for robust performance across diverse downstream tasks.
Our main contributions in this paper are as follows:
  • Introducing a per-head, input-conditioned gating mechanism that dynamically fuses the two parallel attention streams, enabling fine-grained, token-wise noise suppression,
  • Adopting the lateral-inhibition principles of biological systems’ function into artificial attention mechanisms to provide an interpretable contrast-enhancement framework for self-attention,
  • Validating M-DGSA across vision and language benchmarks, where it consistently boosts robustness and cross-domain adaptability under clean and corrupted conditions.
2 Related Work
Contrast-Enhancement and Inhibitory Motifs in Learning.
In biological sensory neurons, lateral inhibition enhances spatial contrast by subtracting pooled neighboring activity, thereby sharpening feature edges and improving signal-to-noise ratio (Hartline, 1940; Kuffler, 1953).
Inspirations from lateral inhibition have informed modern deep networks. These include Local Response Normalization (LRN) which applies fixed local competition to sharpen feature maps (Krizhevsky et al., 2012); and spatial masking methods such as DropBlock, which attenuate contiguous activations to improve generalization (Ghiasi et al., 2018). More recently, LI-CNN learns per-channel subtractive inhibition filters to dynamically emphasize salient patterns (Zhuang et al., 2023), Selective Kernel Networks (SKNet) employ channel-wise gating to adaptively fuse multiple convolutional kernels and dampen irrelevant features, mirroring lateral inhibition (Li et al., 2019), while Gradient Mask applies inhibitory gating on gradients during backpropagation to filter gradient noise and stabilize training (Jiang et al., 2022).
Attention Variants Towards Robustness.
Building on the Transformer (Vaswani et al., 2017) architecture, many attention variants have been proposed to improve efficiency and robustness. Low-rank methods such as Linformer (Wang et al., 2020) and Performer (Choromanski et al., 2020) reduce complexity via kernel approximations, while sparsity-driven models like BigBird (Zaheer et al., 2020) leverage block-sparse patterns. In vision, the Vision Transformer (Dosovitskiy et al., 2020) demonstrates that pure self-attention over image patches, with an extra class token, can rival convolutional networks. ConViT (d’Ascoli et al., 2021) then introduces locality-sensitive gating to blend convolutional inductive biases with self-attention, enabling dynamic control over spatial focus. Recently, the Differential Transformer (Ye et al., 2024) introduces head-wise subtraction of parallel softmax maps to cancel common-mode noise, improve spectral balance and reduce attention collapse.
Despite this rich history, to the best of our knowledge, head-wise, input-dependent lateral-inhibition remains unexplored in multi-head self-attention.
3 Preliminaries
3.1 Scaled Dot-Product Attention
In the original Transformer formulation (Vaswani et al., 2017), given an input sequence of token embeddings
$X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{N\times d_{\mathrm{model}}},$
we compute query, key, and value matrices via learnable linear projections:
$Q=XW_{Q},\quad K=XW_{K},\quad V=XW_{V},$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times d}$. The standard scaled dot-product attention then forms the weight matrix
$A=\mathrm{softmax}\!\bigl{(}\tfrac{QK^{\top}}{\sqrt{d}}\bigr{)}\;\in\mathbb{R}^{N\times N},$
applying the softmax row-wise to normalize each query’s affinities, and the final output is
$\mathrm{Attention}(Q,K,V)=A\,V\;\in\mathbb{R}^{N\times d}.$
3.2 Differential Attention
(Ye et al., 2024) mitigate attention noise by contrasting two parallel attention maps. For an input sequence
$X=[x_{1},\dots,x_{N}]\;\in\;\mathbb{R}^{N\times d_{\mathrm{model}}},$
they first split queries and keys into paired streams and project values jointly:
$[Q_{1};Q_{2}]=X\,W_{Q},\quad[K_{1};K_{2}]=X\,W_{K},\quad V=X\,W_{V},$
with $W_{Q},W_{K}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$ and $W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$. Each stream then forms a standard scaled dot-product attention,
$A_{i}=\mathrm{softmax}\!\bigl{(}\tfrac{Q_{i}K_{i}^{\top}}{\sqrt{d}}\bigr{)}\;\in\mathbb{R}^{N\times N},,\quad i\in\{1,2\},$
and their outputs are combined via a learnable subtraction:
$\mathrm{DiffAttn}(X)=(A_{1}-\lambda\,A_{2})\,V,\in\mathbb{R}^{N\times 2d^{\prime}}.$
Here, $\lambda$ is reparameterized to stabilize learning:
$\lambda=\exp\bigl{(}\lambda_{q1}\lambda_{k1}\bigr{)}-\exp\bigl{(}\lambda_{q2}\lambda_{k2}\bigr{)}+\lambda_{\mathrm{init}},$
where $\lambda_{q1},\lambda_{q2},\lambda_{k1},\lambda_{k2}\in\mathbb{R}^{d}$ are learnable vectors and $\lambda_{\mathrm{init}}=0.8-0.6\,e^{-0.3(l-1)}$ is a layer-dependent constant initialization for layer index $l\in[1,L]$.
3.3 Lateral Inhibition
Lateral inhibition is a fundamental neural mechanism that enhances contrast by suppressing nearby activity. In rate-based models, the neuron’s response is characterized by
$r_{i}=\phi\Bigl{(}e_{i}-\alpha\sum_{j\in\mathcal{N}(i)}e_{j}\Bigr{)},$
where $e_{i}$ is the direct excitatory input, $\sum_{j}e_{j}$ aggregates neighboring signals, $\alpha>0$ controls the inhibition strength, and $\phi$ is a smooth nonlinearity. This local subtractive process accentuates salient features while damping background noise, providing the theoretical underpinning for input-dependent gating in self-attention.
4 Method
We propose Differential Gated Self-Attention (M-DGSA), the first self-attention variant to embed an input-dependent lateral inhibition mechanism, combining the common-mode noise cancellation of differential attention with the adaptability of biological contrast-enhancement. Building on the Differential Transformer (Ye et al., 2024) implementation (dual-softmax streams, GroupNorm (Wu and He, 2018), and $\lambda_{init}$ scheme), we replace its subtraction scalar with a lightweight gating network that learns input-dependent and head-specific inhibitory weights.
4.1 Differential Gated Self-Attention
In our proposed Differential Gated Self-Attention, the input tensor $X\in\mathbb{R}^{N\times d_{model}}$ is linearly projected into two query-key substreams, $Q^{+},K^{+}\in\mathbb{R}^{N\times d^{\prime}}$ (excitatory) and $Q^{-},K^{-}\in\mathbb{R}^{N\times d^{\prime}}$ (inhibitory), alongside a unified value stream $V\in\mathbb{R}^{N\times 2d^{\prime}}$:
$[\,Q^{+};\,Q^{-}\,]=X\,W_{Q},\quad[\,K^{+};\,K^{-}\,]=X\,W_{K},\quad V=X\,W_{V},$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$ are the learnable weight matrices. Then, each substream computes its own scaled dot-product softmax attention map, denoted $A^{+}$ and $A^{-}$:
$A^{+}=\mathrm{softmax}\bigl{(}\tfrac{Q^{+}{K^{+}}^{\top}}{\sqrt{d}}\bigr{)},\quad A^{-}=\mathrm{softmax}\bigl{(}\tfrac{Q^{-}{K^{-}}^{\top}}{\sqrt{d}}\bigr{)}.$
Here we introduce a gating network, a per-input (token) $t$, per-head $h$ sigmoid gate
$g_{t,head_{i}}=\sigma\bigl{(}w_{g,head_{i}}x_{t}+b_{g,head_{i}}\bigr{)},$
which adaptively fuses excitation and inhibition streams
$A_{t,head_{i}}=g_{t,head_{i}}\,A^{+}_{t,head_{i}}\;-\;(1-g_{t,head_{i}})\,A^{-}_{t,head_{i}}.$ (1)
We emphasize subtracting the two gated-weighted terms, rather than blending them via weighted addition as in conventional gating mechanisms. By doing so, each position can dynamically choose how much of the inhibitory branch to apply, amplifying strong signals where needed and quelling noise elsewhere.
Algorithm 1 Multihead Differential Gated Self-Attention (M-DGSA)
1:$X\in\mathbb{R}^{N\times d_{model}}$, layer index $l$, weights $W_{Q},W_{K},W_{V},W_{g},W_{o}$, init $\lambda_{\mathrm{init}}$
2:$\lambda_{l}\leftarrow 0.8-0.6\exp(-0.3(l-1))$
3:$Q\leftarrow XW_{Q},\quad K\leftarrow XW_{K},\quad V\leftarrow XW_{V}$
4:Split $Q$ into $[Q^{+};Q^{-}]$, split $K$ into $[K^{+};K^{-}]$
5:$A^{+}\leftarrow\mathrm{softmax}(Q^{+}K^{+^{\top}}/{\sqrt{d}^{\prime}})$
6:$A^{-}\leftarrow\mathrm{softmax}(Q^{-}K^{-^{\top}}/\sqrt{d}^{\prime})$
7:$g\leftarrow\sigma(XW_{g}+b_{g})$
8:$A\leftarrow g\odot A^{+}-(1-g)\odot A^{-}$
9:$H\leftarrow(1-\lambda_{l})\mathrm{GroupNorm}(AV)$
10:$O\leftarrow$ reshape and concat heads of $H$
11:$Y\leftarrow Q+OW_{o}$
12:$Y\in\mathbb{R}^{N\times hd^{\prime}}$
Figure 1: Design of Multihead Differential Gated Self-Attention. The input tensor is projected into excitatory $(Q^{+},K^{+})$ and inhibitory $(Q^{-},K^{-})$ query-key pairs and shared values $V$. Each branch generates its own softmax map $A^{+}$ and $A^{-}$, which are then "filtered" by a lightweight input-dependent headwise gating network that learns to attenuate noise - mimicking lateral inhibition in biological sensory systems. The fused attention is applied to $V$ and each head’s output undergoes head-wise GroupNorm (Wu and He, 2018) before all heads are concatenated and linearly projected. An optional residual connection around the attention block can be added to improve gradient flow and training stability.
4.2 Multi-head Differential Gated Self-Attention
We extend the DGSA to $h$ heads, as in the original Transformer multi-head attention mechanism (Vaswani et al., 2017). Thereby, we tile the projection matrices into
$[\mathbf{Q^{+}};\mathbf{Q^{-}}],\,[\mathbf{K^{+}};\mathbf{K^{-}}]\;\in\;\mathbb{R}^{N\times 2h\times d^{\prime}},\quad\mathbf{V}\;\in\;\mathbb{R}^{N\times h\times 2d^{\prime}},$
and then reshape into $\{Q^{\pm}_{head_{i}},K^{\pm}_{head_{i}},V_{head_{i}}\}_{1}^{h}$, each of shape $N\times 2\times d^{\prime}$. Each head independently computes its fused map $A_{head_{i}}$ as in Eq. (1) and used to attend over $\mathbf{V}$, in the form of
$H_{head_{i}}=A_{head_{i}}\,V_{head_{i}}.$
This is followed by a GroupNorm (Wu and He, 2018) where RMSNorm (Zhang and Sennrich, 2019) is applied in each head and scaled by $(1-\lambda_{init})$.
Concatenating $[head_{1},\dots,head_{i},\dots,head_{h}]$ along the embedding dimension and projecting with $W_{O}\in\mathbb{R}^{hd^{\prime}\times hd^{\prime}}$ and adding an optional residual connection from $[\mathbf{Q^{+}};\mathbf{Q^{-}}]$ to preserve information flow in our model. This completes the design of our multi-head attention block.
4.3 DGT and DGViT: Transformer & Vision Transformer Instantiations
Our M-DGSA can be applied in any multi-head attention within Transformer-based models. We explore two instantiations:
Differential Gated Transformer (DGT).
Starting from the standard Transformer encoder, we replace every multi-head self-attention block with M-DGSA. Each M-DGSA output is followed by a feed-forward network using the SwiGLU activation (Shazeer, 2020). Unlike in vision tasks, we found that an additional skip connection (shown in Fig. 1) around the M-DGSA block did not improve convergence so we omit it in DGT.
Differential Gated Vision Transformer (DGViT).
Building on the ViT backbone, DGViT replaces each attention module with an M-DGSA layer, retaining the learnable class token in the attention computation and substitutes the GeLU activation (Hendrycks and Gimpel, 2016) in the feed-forward network with SwiGLU. It also integrates a residual connection inside the M-DGSA block to facilitate gradient flow, stabilize training, and boost model’s performance.
5 Experiments
In this section, we consider two important domains – vision and language classification - where we evaluate the performance of our differential gated self-attention mechanism. Detailed experimental setup and hyperparameter choices are provided in the supplementary materials.
All models are trained from scratch (no pre-trained weights), so our absolute accuracies may be lower than those reported in studies that fine-tune pre-trained networks. All experiments are implemented in PyTorch 2.7.0 and ran on a single machine equipped with an NVIDIA A10G GPU (24 GB VRAM).
5.1 Vision Classification
For vision classification, we leverage the built-in torchvision.datasets module, which offers a wide range of standardized benchmarks that integrate seamlessly with PyTorch for efficient data loading and preprocessing transformations. To evaluate robustness under challenging conditions, we inject noise into the training images across multiple distortion levels and noise types.
Datasets.
CIFAR-10 (Krizhevsky et al., 2009) contains 32×32 color images evenly distributed across 10 classes, and CIFAR-100 is its 100-class counterpart. Fashion-MNIST (Xiao et al., 2017) comprises 28×28 pixel grayscale Zalando article images spanning 10 fashion categories. The Street View House Numbers (SVHN) dataset (Netzer et al., 2011) consists of real-world house number images extracted from Google Street View.
Models.
We compare our DGViT model against the original Vision Transformer (ViT) architecture, in which the feed-forward module expands the hidden dimension by a factor of two or four, applies a SwiGLU activation, and then projects back to the original size. To examine the impact of regularization, we insert one or two dropout layers interleaved between these projection and activation steps, thus assessing performance under different levels of stochastic masking throughout training.
Results.
Table 1 shows that the optimal configuration uses a 4×4$\times$4 × expansion followed by SwiGLU activation and a projection back to $d$, with two dropout layers. Relative to the baseline ViT, our DGViT model delivers competitive performance and marginal gains across all evaluated image-classification benchmarks. Notably, on CIFAR-10 we observe an absolute accuracy increase exceeding 2%. These findings indicate that expanding-and-shrinking the dimensionality around a SwiGLU activation, together with strategically placed dropout layers, promotes robust generalization across diverse visual domains and downstream tasks.
Table 1: Classification accuracy (mean $\pm$ std) on vision benchmarks. Averaged over 5 seeds.
CIFAR-10 (%) CIFAR-100 (%) FashionMNIST (%) SVHN (%)
#Classes 10 100 10 10
#Size 60,000 60,000 70,000 600,000
ViT $75.90\pm 0.09$ $47.01\pm 0.56$ $93.15\pm 0.11$ $93.22\pm 0.28$
DGViT (ours)
   proj. drop.
   ×$\times 2$$\times 2$ $77.51\pm 0.44$ $47.13\pm 0.26$ $93.42\pm 0.07$ $93.93\pm 0.34$
   ×$\times 4$$\times 1$ $77.77\pm 0.29$ $47.32\pm 0.38$ $93.32\pm 0.1$ $\mathbf{94.16\pm 0.49}$
   ×$\times 4$$\times 2$ $\mathbf{78.06\pm 0.41}$ $\mathbf{47.72\pm 0.35}$ $\mathbf{93.45\pm 0.13}$ $93.94\pm 0.18$
5.2 Language Classification
In our language classification experiments, we progressively evaluate our model on tasks of increasing difficulty by the number of target classes. We begin with binary classification involving 2 classes, advance to a 4-class setting, and finally test the model’s capacity on a more complex 20-class classification task. This gradual increase in the number of classes allows us to systematically assess the model’s performance and robustness across varying levels of classification complexity.
Datasets.
Each dataset comes with pre-split training and test sets and we derive a validation split by dividing the original training set into 80% for training and 20% for validation. The Rotten Tomatoes  (Pang and Lee, 2005) and the IMDB  (Maas et al., 2011) movie reviews datasets are for binary sentiment classification. The AG News dataset (Del Corso et al., 2005), derives from AG’s corpus of news articles, focused on four major news categories. The 20 Newsgroups dataset (Mitchell, 1997) contains newsgroup messages organized into 20 distinct categories.
Table 2: Classification accuracy (mean ± std) on language benchmarks. Averaged over 5 seeds.
Rotten Tomatoes (%) IMDB (%) AGNews (%) 20 Newsgroups (%)
#Classes 2 2 4 20
#Size 10,000 50,000 127,600 18,000
Transformer $72.18\pm 0.51$ $85.34\pm 0.22$ $91.61\pm 0.08$ $51.54\pm 0.49$
DIFF Transformer
   proj.
   ×$\times 2$ $73.99\pm 0.91$ $86.41\pm 0.14$ $91.85\pm 0.23$ $46.38\pm 0.96$
   ×$\times 4$ $73.06\pm 0.7$ $86.49\pm 0.29$ $92.03\pm 0.2$ $47.38\pm 2.08$
   ×$\times 16/3$ $73.19\pm 0.85$ $86.43\pm 0.16$ $92.00\pm 0.19$ $49.99\pm 2.80$
DGT (ours)
   proj.
   ×$\times 2$ $74.26\pm 0.97$ $86.47\pm 0.17$ $\mathbf{92.25\pm 0.14}$ $60.32\pm 0.49$
   ×$\times 4$ $73.83\pm 1.15$ $\mathbf{86.97\pm 0.43}$ $91.89\pm 0.14$ $59.07\pm 0.48$
   ×$\times 16/3$ $\mathbf{74.40\pm 0.42}$ $86.53\pm 0.33$ $91.98\pm 0.27$ $\mathbf{63.53\pm 0.87}$
Models.
We compare our DGT model’s performance to both the vanilla Transformer and the Differential Transformer under default configurations. For a fair comparison, we evaluate Differential Transformer variants with feed-forward modules expanded by factors of 2, 4, and 16/3 relative to the output dimension $Y$, applying SwiGLU before projecting back to the base dimension, where the 16/3 expansion matches the original Differential Transformer setup. We then apply these identical expansion factors and activation sequence to our DGT model, aligning experimental conditions across all architectures to isolate the impact of our approach across multiple benchmark tasks and comprehensive evaluation protocols.
Results.
Our results in Table 2 indicate that incorporating Differential Gated Attention improves prediction accuracy in language classification tasks. Specifically, we observe that the DGT model’s performance increases as the complexity of the classification task grows, particularly with a higher number of target classes. While the model achieves a modest improvement of 1-2% on the Rotten Tomatoes, IMDB, and AG News datasets, it shows a substantial gain of 12-14% on the more challenging 20 Newsgroups task. This trend suggests that the model not only scales well with increased task difficulty but also benefits from the richer structure and diversity present in fine-grained classification settings.
5.3 Attention Visualization.
We use the attention-rollout method (Abnar and Zuidema, 2020) to inspect how our models filter noise and sharpen structure on representative vision and language examples.
Figure 2: Attention-rollout heatmaps for DGViT vs ViT on CIFAR-100 (top two rows) and SVHN (bottom two rows). Bright yellow indicates high attention weights, whereas deep blue denotes low or suppressed attention. While ViT’s attention is broadly distributed, assigning high weights to both the target and surrounding background pixels, DGViT exhibits greater discriminative focus.
Vision (DGViT vs ViT).
Fig. 2 shows attention-rollout heatmaps for DGViT versus the standard ViT on CIFAR-100 and SVHN. In each row, the left panel is the input image, the middle, DGViT’s fused attention, and the right, ViT’s attention. DGViT concentrates excitatory weights (bright yellow) on semantically salient structures - e.g., mushroom contours and digit strokes - while assigning inhibitory weights (deep blue) to background clutter, highlighting its fine-grained boundary delineation and illustrating selective focus on semantically salient regions.
(a) Transformer model attention
(b) DIFF-Transformer model attention
(c) DGT (ours) model attention
Figure 3: Attention-rollout for a 20 Newsgroups sample. Bright yellow indicates strong positive attention, and deep blue denotes inhibitory weights. (a) The standard Transformer baseline spreads its attention diffusely across both function words and content tokens, failing to isolate the class-defining terms. (b) The Differential Transformer partially suppresses common-mode noise, but still assigns non-negligible weight to other classes too. (c) Our DGT model sharply concentrates on the key parts of the input while attenuating filler words. The focused, input-dependent gating aligns attention with the correct comp.windows.x label and filters out background noise.
Language (DGT vs Baselines).
Fig. 3 presents attention-rollout for (a) Transformer, (b) Differential Transformer, and (c) our DGT on a 20 Newsgroups sample, input text: "Most graphics systems I have seen have drawing routines that also specify a color for drawing, like Drawpoint(x,y,color) or Drawline(x1,y1,x2,y2,color) or Fillrectangle(x1,y1,x2,y2,color)…With X, I…", label: comp.windows.x. The vanilla Transformer disperses its weights broadly over both common function words and various candidate labels, failing to single out the true class while the Differential Transformer partly suppresses common-mode noise yet still spreads attention to competing classes, diluting its focus on the correct one. DGT’s input-dependent gating, in contrast, concentrates focus on key class-indicative words, filtering out irrelevant text and yielding a more interpretable, discriminative attention pattern, increasing attention mass on ground-truth and aligning sharply with comp.windows.x.
6 Discussion
Our results demonstrate that embedding an input-conditioned lateral inhibition mechanism into self-attention yields substantial gains in noise resilience and generalization across both vision and language tasks. M-DGSA consistently sharpens attention maps, assigning strong excitatory weights to semantically relevant regions while suppressing background noise, without requiring any task-specific tuning. Its capacity to delineate fine boundaries suggests applications in safety-critical domains - such as lane-marking detection in autonomous driving or edge delineation in medical imaging—where robust, fine-grained attention is paramount.
Despite these benefits, several limitations invite further investigation. First, we have evaluated on mid-scale benchmarks with synthetic noise, extending to real-world distortions (e.g., motion blur, occlusions) and large-scale tasks (e.g., ImageNet classification) may reveal additional strengths or failure modes. Second, we confined M-DGSA to self-attention, integrating lateral inhibition into cross-attention, such as in vision-language models for VQA or image captioning, could enable dynamic, question-conditioned suppression of irrelevant image regions, further enhancing multimodal alignment.
For future work, we see several promising directions: extending lateral-inhibition gating to cross-attention and multimodal models, scaling to large-scale benchmarks and real-world corruption regimes, and developing hardware-efficient implementations that exploit the induced attention sparsity. We also anticipate that a deeper theoretical understanding of learned inhibition patterns will inform principled design and initialization of future attention mechanisms.
7 Conclusion
We have introduced Multihead Differential Gated Self-Attention (M-DGSA), a lightweight yet powerful extension of multihead self-attention that embeds per-head input-dependent lateral inhibition into the dual-softmax framework. Leveraging dynamic, context-sensitive gating, M-DGSA achieves sharper, more selective attention maps, delivering enhanced noise resilience and improved cross-domain performance on both vision and language benchmarks. Importantly, it integrates seamlessly into existing Transformer and ViT architectures with negligible overhead, making it immediately applicable to a wide range of tasks.
Beyond empirical improvements, M-DGSA advances the interpretability of attention mechanisms: its gating patterns mirror biological contrast-enhancement, providing clear, roll-out visualizations of how models amplify salient features and quench noise. We believe this biologically inspired paradigm marks a significant step toward more robust, trustworthy attention models for next-generation deep learning.
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034277.
Appendix A Additional Experimental Details
A.1 Detailed description of the used datasets
Rotten Tomatoes. (Pang and Lee, 2005) $10,662$ movie reviews equally split between positive and negative sentiment.
IMDB. (Maas et al., 2011) $50,000$ highly polarized reviews ($25,000$ train / $25,000$ test) for binary sentiment classification.
AG News. (Del Corso et al., 2005) Derived from AG’s corpus of news articles, the AG News dataset comprises four categories—World, Sports, Business, and Science/Technology—by combining each article’s title and description and selecting the four largest classes. It contains $120,000$ training samples ($30,000$ per category) and $7,600$ test samples ($1,900$ per category).
20 Newsgroups. (Mitchell, 1997) $\sim 18,000$ posts organized into 20 distinct newsgroups, each corresponding to a specific topic such as computer hardware, recreational activities, politics, science, or religion. We adopt the canonical split of $11,314$ training and $7,532$ test posts, yielding on average about $566$ training and $377$ test samples per class.
CIFAR-10 / CIFAR-100. (Krizhevsky et al., 2009) $60,000$ colored images of size 32x32 pixels, evenly distributed across 10 distinct classes such as airplanes, automobiles, birds, cats, deers, dogs, frogs, horses, ships, and trucks. Each class includes $6,000$ images, with the dataset split into $50,000$ training and $10,000$ test images. CIFAR-100 is based on the same concept but with 100 classes and contains 600 samples for each class.
FashionMNIST. (Xiao et al., 2017) $70,000$ 28×28 pixel grayscale Zalando article images of 10 fashion item categories (e.g., t-shirt/top, trouser, bag, sandal). Maintaining the same structure and format as MNIST, FashionMNIST provides a standardized yet more complex alternative for evaluating machine learning models on image classification tasks.
Street View House Numbers (SVHN). (Netzer et al., 2011) A large-scale, real-world image dataset comprising over 600,000 32×32-pixel digit crops extracted from Google Street View house numbers. It includes 10 classes (digits 0–9): 73,257 training images, 26,032 test images, and 531,131 additional “extra” samples of comparatively easier digits. By capturing digits in varied natural scenes—with diverse backgrounds, lighting conditions, and occlusions—SVHN poses a substantially more challenging recognition task than MNIST.
Licenses for existing assets
A.2 Dataset Splits
For each language dataset, AG News, IMDB, 20 Newsgroups, and Rotten Tomatoes, we further split the training fold into $80\%$ train and $20\%$ validation (stratified by class). All final results are reported on the official test set. For all vision benchmarks, CIFAR-10, CIFAR-100, MNIST, Fashion-MNIST, and SVHN, we use the standard train/test partitions provided by each dataset and no further hold-out split was applied.
A.3 Hyperparameters
A.3.1 Vision Benchmarks
Table 3: Hyperparameters for vision datasets.
Hyperparameter Value
Epochs 100
Batch size 128
Learning rate $3\times 10^{-4}$
Embedding dimension 256
# Layers 8
# Heads 8
Dropout probability 0.05
Weight decay 0.01
Optimizer AdamW ($\beta_{1}=0.9,\;\beta_{2}=0.999,\;\epsilon=10^{-8}$)
LR scheduler CosineAnnealingLR
Patch size 4
Max sequence length 100
A.3.2 Language Benchmarks
All text-classification experiments use the hyperparameters listed in Table 4. For the Rotten Tomatoes dataset, we instead employ 4 layers, a learning rate of $45e-$ 4, and 500 warmup steps.
Table 4: Hyperparameters for text datasets.
Hyperparameter Value
Epochs 10
Batch size 32
Learning rate $3\times 10^{-4}$
Warmup steps 1000
Max sequence length 256
Minimum token frequency 2
Maximum vocabulary size 60,000
Embedding dimension 256
# layers 6
# heads 8
Dropout probability 0.1
Optimizer AdamW ($\beta_{1}=0.9,\;\beta_{2}=0.98,\;\epsilon\!=\!10^{-8}$)
Weight decay 0.01 (0 for bias & LayerNorm/RMSNorm weights)
LR scheduler Linear warmup (1000 steps) + linear decay
Gradient clipping Max-norm = 1.0
A.4 Compute & Runtime
All experiments ran in PyTorch on a single NVIDIA A10G (24 GB VRAM).
  • Vision: $\sim 0.5$min/epoch on CIFAR-10/100, $\sim 0.6$ min on FashionMNIST, $\sim 0.76$ min on SVHN.
  • Language: $\sim 0.3$min/epoch on Rotten Tomatoes, $\sim 0.39$ min on IMDB, $\sim 0.57$ min on AG News, $\sim 0.5$ min on 20NewsGroups
Table 5: Number of learnable parameters per epoch for vision benchmarks.
CIFAR-10 CIFAR-100 FashionMNIST SVHN
   DGViT
   proj. ×$\times 2$ $3.77$M $3.8$M $3.77$M $3.77$M
   proj. ×$\times 4$ $5.35$M $5.36$M $5.35$M $5.35$M
Table 6: Number of learnable parameters per epoch for language benchmarks.
Rotten Tomatoes IMDB AGNews 20 Newsgroups
   DGT
   proj. ×$\times 2$ $3.8$M $25.7$M $13.2$M $12$M
   proj. ×$\times 4$ $4.6$M $26.9$M $14.4$M $12.8$M
   proj. ×$\times 16/3$ $5.1$M $27.7$M $15.2$M $13.4$M
A.5 Additional Analyses
Ablation on $\lambda_{init}$
After experimenting with various $\lambda_{init}$ settings, we found that holding it fixed at 0.8 outperforms both other constant values and treating it as a learnable parameter.
Gate depth
We observed that increasing the depth of the MLP gating layers consistently degraded the performance, therefore, our method employs a single-layer gating mechanism.